# HOIDiNi: Human-Object Interaction through Diffusion Noise Optimization

### v1

@@ Appendix B Evaluation Metrics
We evaluate our method along two dimensions: motion realism and interaction accuracy. Following IMoS (Ghosh et al., 2023), we compute realism metrics using embeddings from the final layer of an intent classifier. However, the classifier used in IMoS is limited to body joint positions and cannot capture fine-grained grasp dynamics. In contrast, our evaluation employs a more expressive classifier that takes as input body joints, hand joints, and object trajectories, allowing for a more comprehensive assessment of interaction quality.
FID
@@
Fréchet inception distance measures the distance between the distributions of generated and ground-truth motions in a learned embedding space. Lower FID values indicate that the synthetic motion is closer in distribution to real motion data, capturing both realism and diversity.
Diversity
@@
Evaluates how varied the generated motions are across different samples for the same input condition (e.g., prompt or object). It is computed as the average pairwise distance between multiple motion samples in the embedding space. A lower difference between ground truth and generated diversity scores suggests that the generated motions effectively capture the observed variability of human movement.
AVE
@@
Measures the discrepancy between the variance of joint positions in generated motion and that of ground-truth motion. Specifically, it computes the average $L^{2}$ difference in per-joint positional variance across time. A lower AVE suggests that the model accurately captures the temporal dynamics and variability of natural human movement, avoiding overly rigid or overly jittery outputs.
IRA
@@
Intent recognition accuracy quantifies how well the generated motions conveys the intended interaction or action. It is computed as the classification accuracy of the intent classifier on generated samples. High IRA indicates that the generated motions are semantically meaningful and align with their intended action labels, providing a measure of goal consistency and plausibility.
@@
Figure 9. Palm anchor set $\mathcal{A}$ used by CPHOI
Figure 10. A screenshot from the user study.
Multimodality
@@
Assesses the model’s capacity to produce distinct motions for the same conditioning intent. Unlike diversity, which measures sample variation globally, multimodality focuses on conditional variability by comparing multiple outputs conditioned on the same prompt. This metric is crucial for evaluating whether the model can express different plausible interaction strategies.
Penetration
@@
Quantifies physical implausibility by measuring the extent to which the human mesh intersects with the object mesh. We compute the mean maximal penetration depth across frames where interpenetration occurs and the object is above table height. Lower penetration values indicate more physically valid interactions, particularly in grasping and manipulation scenarios where accurate surface contact is essential.
Floating
@@
Captures the failure of hand-object interaction where the hand remains unnaturally far from the object surface. It is computed as the mean shortest distance between the body and object meshes, averaged across all motions (excluding frames with penetration and frames where the object is at table height). High floating values typically reflect unrealistic, disconnected grasping motion.
Appendix C User Study
@@
We conducted a user study with 24 participants, evaluating, in total, 12 side-by-side randomly selected samples of two models using the same inputs. We asked the user to evaluate the grasp quality, prompt adherence, and overall quality. As shown in Figure 7, users preferred the results generated by our framework. A representative screenshot from the study interface is shown in Figure 10.

### v2

@@ D.1 GRAB Evaluation
For the GRAB dataset (Taheri et al., 2020a) experiment, we follow IMoS Ghosh et al. (2023), and compute realism metrics using embeddings from the final layer of an intent classifier. However, the classifier used in IMoS is limited to body joint positions and cannot capture fine-grained grasp dynamics. In contrast, our evaluation employs a more expressive classifier that takes as input body joints, hand joints, and object trajectories, allowing for a more comprehensive assessment of interaction quality.
@@
FID. Fréchet inception distance measures the distance between the distributions of generated and ground-truth motions in a learned embedding space. Lower FID values indicate that the synthetic motion is closer in distribution to real motion data, capturing both realism and diversity.
@@
Diversity. Evaluates how varied the generated motions are across different samples for the same input condition (e.g., prompt or object). It is computed as the average pairwise distance between multiple motion samples in the embedding space. A lower difference between ground truth and generated diversity scores suggests that the generated motions effectively capture the observed variability of human movement.
@@
AVE. Measures the discrepancy between the variance of joint positions in generated motion and that of ground-truth motion. Specifically, it computes the average $L^{2}$ difference in per-joint positional variance across time. A lower AVE suggests that the model accurately captures the temporal dynamics and variability of natural human movement, avoiding overly rigid or overly jittery outputs.
@@
IRA. Intent recognition accuracy quantifies how well the generated motions conveys the intended interaction or action. It is computed as the classification accuracy of the intent classifier on generated samples. High IRA indicates that the generated motions are semantically meaningful and align with their intended action labels, providing a measure of goal consistency and plausibility.
Figure 9: A screenshot from the user study.
@@
Figure 10: Palm anchor set $\mathcal{A}$ used by CPHOI for the GRAB (Taheri et al., 2020a) experiment.
@@
Multimodality. Assesses the model’s capacity to produce distinct motions for the same conditioning intent. Unlike diversity, which measures sample variation globally, multimodality focuses on conditional variability by comparing multiple outputs conditioned on the same prompt. This metric is crucial for evaluating whether the model can express different plausible interaction strategies.
@@
Penetration. Quantifies physical implausibility by measuring the extent to which the human mesh intersects with the object mesh. We compute the mean maximal penetration depth across frames where interpenetration occurs and the object is above table height. Lower penetration values indicate more physically valid interactions, particularly in grasping and manipulation scenarios where accurate surface contact is essential.
@@
Floating. Captures the failure of hand-object interaction where the hand remains unnaturally far from the object surface. It is computed as the mean shortest distance between the body and object meshes, averaged across all motions (excluding frames with penetration and frames where the object is at table height). High floating values typically reflect unrealistic, disconnected grasping motion.
D.2 OMOMO Evaluation
For the OMOMO dataset (Li et al., 2023) experiment, we follow the evaluation metrics defined by CHOIS (Li et al., 2024a):
Condition Matching Metric.
Those metrics calculate the Euclidean distance between the predicted and input object waypoints. It includes the start and end position errors $\left(T_{s},T_{e}\right)$, and waypoint errors $\left(T_{xy}\right)$ measured in centimeters (cm).
Human Motion Quality Metric.
Those metrics encompasses the foot sliding score (FS), foot height $\left(H_{feet}\right)$, Fréchet Inception Distance $\left(FID\right)$ and R-precision $\left(R_{prec}\right)$. FS is the weighted average of accumulated translation in the xy plane, following prior work He et al. (2022), measured in centimeters (cm). $H_{feet}$ assesses the height of the feet, also in centimeters. $R_{prec}$ and $FID$ are computed following the text-to-motion task Guo et al. (2022a). $R_{prec}$ (top-3) measures whether the generated motion is consistent with the text. $FID$ assesses the motion quality by computing the discrepancy between the distributions of ground truth and generated motions.
Interaction Quality Metrics.
Those metrics assess the accuracy of hand-object interactions, encompassing both contacts and penetrations. For contact accuracy, it employs precision $\left(C_{prec}\right)$, recall $\left(C_{rec}\right)$, and F1 score $\left(C_{F_{1}}\right)$ metrics following prior work Li et al. (2023). Additionally, it includes contact percentage $\left(C_{\%}\right)$, determined by the proportion of frames where contact is detected. To compute the penetration score $\left(P_{hand}\right)$, each vertex of the hand $V_{i}$ is used to query the precomputed object's Signed Distance Field (SDF). This process yields a corresponding distance value $d_{i}$ for each vertex. The penetration score is then derived by computing the average of the negative distance values (representing penetration), formalized as $\frac{1}{n}\sum_{i=1}^{n}|min(d_{i},0)|$, measured in centimeters (cm).
We note that CHOIS additionally measured the distance of the generated motion from the corresponding ground truth motion. Since HOIDiNi is a generative model, not aiming to reconstruct the ground truth, we find this metric irrelevant for our scope and omit it.
Appendix E User Study
@@
We conducted a user study for the GRAB dataset (Taheri et al., 2020a) with 24 participants, evaluating, in total, 12 side-by-side randomly selected samples of two models using the same inputs. We asked the user to evaluate the grasp quality, prompt adherence, and overall quality. As shown in Figure 7, users preferred the results generated by our framework. A representative screenshot from the study interface is shown in Figure 10.
Appendix F LLM Usage
In this paper, we used ChatGPT 4o/5 to revise our writing, code assisting, and to enhance the GRAB (Taheri et al., 2020a) labels into text prompts.